# Joint Activity Detection and Channel Estimation for Massive Connectivity: Where Message Passing Meets Score-Based Generative Priors

### v1

@@ V-B1 Damping
The pseudo AWGN observation model in (20) is justified only in the large system limit with right-rotationally invariant random $\mathbf{Q}$ [50]. In practice, the equivalent noise may not be strictly Gaussian due to the particular choice of $\mathbf{Q}$ and the finite dimensions of $M$, $N$, $K$, and $T$. This discrepancy can lead to some unexpected numerical issues and diverges. In this work, we employ the damping technique to improve the numerical stability of STMP-JADCE. Specifically, in each iteration, we smoothen the updates of the means and variances by using a linear combination of the current and previous updates. For example, we replace the input to module A (Line 15 of Algorithm 1) at the $i$-th iteration by
@@
V-B2 Channel Power Normalization
Similarly, the scaled input of $\tau^{\mathtt{pri}}$ is given by
@@ V-B2 Channel Power Normalization
V-C Pilot Design
@@ VI-A1 Channel Dataset
We adopt the CDL family of channel models generated from the QuaDRiGa toolbox [53] for training and evaluation. The training and evaluation sets consist of $50,000$ and $5,000$ channel samples, respectively. The evaluation samples are directly generated from the CDL-C model specified in [31, Table 7.7.1-3]. In the training set, we randomly generate the angles, delays, and complex gains of each path in CDL-C to cover a diverse range of channel realizations. The CDL-C channel model comprises $24$ clusters, each contributing $20$ sub-paths, resulting in a total of $L=480$ propagation paths.
@@
In simulation, the BS is equipped with a uniform linear array (ULA) consisting of $M=32$ antennas, arranged with half-wavelength spacing. A total of $N=48$ OFDM subcarriers is considered, operating in a central frequency of $2.6$ GHz. The subcarrier spacing is $\Delta f=30$ kHz. The delay spread is randomly chosen in the range of $[100,363]$ ns. The channel model can be expressed as $\mathbf{H}=\sum_{l=1}^{L}\beta_{l}\sqrt{\rho_{l}}\mathbf{a}_{N}(\tau_{l})\mathbf{b}_{M}^{\sf H}(\varphi_{l})$, where $\beta_{l}$, $\rho_{l}$, $\tau_{l}$, and $\varphi_{l}$ denote the complex coefficient, the path power, the path delay, and the angle of arrival (AoA) of the $l$-th path, respectively; the temporal and spatial steering vectors are defined as $\mathbf{a}_{N}(\tau_{l})\triangleq\frac{1}{\sqrt{N}}\left[1,e^{-j2\pi\Delta f\tau_{l}},\dots,e^{-j2\pi(N-1)\Delta f\tau_{l}}\right]^{\sf T}$ and $\mathbf{b}_{M}(\varphi_{l})\triangleq\frac{1}{\sqrt{M}}\left[1,e^{-j\pi\sin\varphi_{l}},\dots,e^{-j\pi(M-1)\sin\varphi_{l}}\right]^{\sf T}$. The distance between the devices and the BS ranges from $35$ to $200$ m. The path-loss component is modeled as $-128.1-36.7\log_{10}d_{k}$ in dB scale, where $d_{k}$ is the distance (in km) between the $k$-th device and the BS.
VI-A2 Training Configurations
@@
We adopt the NCSNv2 [54] and the NCSN [15] networks as the architectures for the first- and the second-order score models, respectively. The output of NCSN is the diagonals of the second-order score function, yielding the same dimension as the first-order score function. The models are trained on every integer SNR in the range of $[-10,30]$ dB, resulting in a total of $I=41$ distinct SNR values. The noise levels $\{\tau_{i}\}_{i=1}^{I}$ during training are set accordingly. We use the Adam optimizer for both models. We select a batch size of $400$ and a learning rate of $0.0001$. The number of training epochs is set to $400$. Moreover, the exponential moving average (EMA) of model weights [54] is applied to enhance the training stability, in which the EMA rate is chosen as $0.999$. The training is conducted on a Linux server with 4 NVIDIA A800 GPUs and an Intel Xeon Platinum 8358P CPU @ 2.60 GHz.
@@ VI-A3 Evaluation Settings
During evaluation, the channel of each device is randomly sampled from the evaluation set. The following settings are adopted unless specified otherwise: $\lambda=0.05$, $K=800$, and the SNR is $10$ dB. For the figures showing the NMSE performance and detection error probability, each data point is averaged over $500$ and $5000$ Monte Carlo trials, respectively. The evaluation utilizes a single NVIDIA A800 GPU, which is performed on the same machine used for training.
@@ VI-C JADCE Performance
Fig. 6 shows the NMSE in channel estimation versus the SNR. Our method demonstrates a significant improvement over the baselines across all SNRs. The performance gap between our method and BG-TMP highlights the critical importance of an accurate channel prior in message passing. Moreover, the TMP-EM algorithm favors a smaller number of sub-blocks at low SNR, while preferring a larger number of sub-blocks at high SNR. The sensitivity to the number of sub-blocks can pose challenges in hyperparameter tuning.
In Fig. 7, we present the detection error probability, which sums up missed detections and false alarms, plotted against different SNRs. The BG-TMP algorithm often fails to converge and is not included. Interestingly, although all schemes utilize the same Bernoulli prior for device activity, our method achieves the lowest detection error probability. This improvement can also be attributed to the plug-in of the score-based generative channel prior. The reason is the following. As shown in Fig. 6, the accurate channel prior brings a remarkable gain in channel estimation. Through message passing iterations, the device activity is detected using this more accurate channel estimate, hence resulting in enhanced detection performance.
@@
In Fig. 8, we plot the achieved NMSE in channel estimation for various numbers of total devices. Targeted at a channel estimation NMSE of $-25$ dB, our method can support $3000$ connected devices, while the best baseline (TMP-EM with $12$ sub-blocks) can accommodate only $750$ devices, leading to a 4×4$\times$4 × increase in access capacity. Moreover, for NMSE values below $-5$ dB, STMP-JADCE shows the flattest slope in the performance curve. This suggests that, given an equivalent tolerance in NMSE degradation (in dB), our approach can support the largest increase in the number of connected devices. These observations highlight the superiority of the proposed algorithm for achieving massive connectivity.
@@ VII Conclusion and Discussion
In this paper, we introduced STMP-JADCE for massive connectivity, a model-based deep learning solution that combines message passing with score-based generative models. In each message passing iteration, our approach maintains both a first-order and a second-order score network to calculate the posterior mean and variance of the channel, respectively. We show by simulation results that STMP-JADCE drastically outperforms all competing algorithms, while achieving a comparable or even faster convergence speed.

### v2

@@ V-B1 Damping
The pseudo AWGN observation model in (20) is justified only in the large system limit with right-rotationally invariant random $\mathbf{Q}$ [51]. In practice, the equivalent noise may not be strictly Gaussian due to the particular choice of $\mathbf{Q}$ and the finite dimensions of $M$, NN, $K$, and $T$. This discrepancy can lead to some unexpected numerical issues and diverges. In this work, we employ the damping technique to improve the numerical stability of STMP-JADCE. Specifically, in each iteration, we smoothen the updates of the means and variances by using a linear combination of the current and previous updates. For example, we replace the input to module A (Line 15 of Algorithm 1) at the $i$-th iteration by
@@
V-B2 Channel Power Normalization
Similarly, the scaled input of $\tau^{\mathtt{pri}}$ is given by
V-C Pilot Design and Complexity Analysis
@@
The computational complexity of STMP-JADCE is evaluated in terms of the number of complex multiplications. The main contributors are the matrix-vector multiplications in module A and the forward passes through the score networks in module B. In particular, each matrix-vector multiplication in (7) can be implemented by NN length-$K$ FFT/IFFT operations, resulting in a total of $2MNK\log_{2}K$ complex multiplications across all $M$ antennas. Other element-wise operations are of lower order and therefore neglected. The complexity of the score networks is denoted by $C=\max\{C_{1},C_{2}\}$, where $C_{1}$ and $C_{2}$ represent the computational costs of a forward pass through the first- and second-order score networks, respectively. Consequently, the overall per-iteration complexity of STMP-JADCE is $\mathcal{O}\left(MNK\log_{2}K+C\right)$.
@@ VI-A1 Channel Dataset
We adopt the CDL family of channel models generated from the QuaDRiGa toolbox [55] for training and evaluation. The training and evaluation sets consist of $50,000$ and $5,000$ channel samples, respectively. The evaluation samples are directly generated from the CDL-C model specified in [32, Table 7.7.1-3]. In the training set, we randomly generate the angles, delays, and complex gains of each path in CDL-C to cover a diverse range of channel realizations. The CDL-C channel model comprises $24$ clusters, each contributing $20$ sub-paths, resulting in a total of $L=480$ propagation paths.
@@
In simulation, the BS is equipped with a uniform linear array (ULA) consisting of $M=32$ antennas, arranged with half-wavelength spacing. A total of $N=48$ OFDM subcarriers is considered, operating in a central frequency of $2.6$ GHz. The subcarrier spacing is $\Delta f=30$ kHz. The delay spread is randomly chosen in the range of $[100,363]$ ns. The channel model can be expressed as $\mathbf{H}=\sum_{l=1}^{L}\beta_{l}\sqrt{\rho_{l}}\mathbf{a}_{N}(\tau_{l})\mathbf{b}_{M}^{\sf H}(\varphi_{l})$, where $\beta_{l}$, $\rho_{l}$, $\tau_{l}$, and $\varphi_{l}$ denote the complex coefficient, the path power, the path delay, and the angle of arrival (AoA) of the $l$-th path, respectively; the temporal and spatial steering vectors are defined as $\mathbf{a}_{N}(\tau_{l})\triangleq\frac{1}{\sqrt{N}}\left[1,e^{-j2\pi\Delta f\tau_{l}},\dots,e^{-j2\pi(N-1)\Delta f\tau_{l}}\right]^{\sf T}$ and $\mathbf{b}_{M}(\varphi_{l})\triangleq\frac{1}{\sqrt{M}}\left[1,e^{-j\pi\sin\varphi_{l}},\dots,e^{-j\pi(M-1)\sin\varphi_{l}}\right]^{\sf T}$. The distance between the devices and the BS ranges from $35$ to $200$ m. The path-loss component is modeled as $128.1-36.7\log_{10}d_{k}$ in dB scale, where $d_{k}$ is the distance (in km) between the $k$-th device and the BS.
VI-A2 Network Architectures and Hyperparameters
We adopt the NCSNv2 [56] and NCSN [15] architectures for the first- and second-order score networks, respectively. Both networks adopt multi-scale convolutional architectures based on a U-Net or RefineNet backbone, composed of stacked residual and refinement blocks. Each block consists of two $3\times 3$ convolutions followed by instance normalization and nonlinear activation, with skip connections between corresponding encoder and decoder stages to preserve fine-grained spatial information. The networks progressively downsample feature maps from $128$ to $256$ channels through residual blocks, then upsample them symmetrically using refinement blocks with dilated convolutions to enlarge the receptive field. The noise level is embedded through a conditioning mechanism implemented by a small multilayer perceptron followed by feature-wise linear modulation, which injects noise-dependent scale and shift parameters into every block. The output of the NCSN model corresponds to the diagonal entries of the second-order score (Hessian) function, sharing the same dimension as the first-order score. The detailed layer configurations of both architectures are summarized in Table III.
@@ VI-A2 Network Architectures and Hyperparameters
The models are trained on every integer SNR in the range of [10,30][-10,30] dB, yielding $I=41$ distinct noise levels $\{\tau_{i}\}_{i=1}^{I}$. Both networks are optimized using Adam with a learning rate of $10^{-4}$, batch size $400$, and $400$ training epochs. An exponential moving average (EMA) of the model parameters with rate $0.999$ is applied to stabilize training [56]. The training is conducted on a Linux server with 4 NVIDIA A800 GPUs and an Intel Xeon Platinum 8358P CPU @ 2.60 GHz. Under this configuration, the training of the first- and second-order score networks each takes approximately one day.
TABLE III: Network Architectures
First-order score network (NCSNv2) Second-order score network (NCSN)
$\times3$ Conv2D, 128 $\times3$ Conv2D, 128
ResBlock, 128 CondResBlock, 128
ResBlock, 128 CondResBlock, 128
ResBlock down, 256 CondResBlock down, 256
ResBlock, 256 CondResBlock, 256
ResBlock down, 256 (dilation 2) CondResBlock down, 256 (dilation 2)
ResBlock, 256 (dilation 2) CondResBlock, 256 (dilation 2)
ResBlock down, 256 (dilation 4) CondResBlock down, 256 (dilation 4)
ResBlock, 256 (dilation 4) CondResBlock, 256 (dilation 4)
RefineBlock, 256 CondRefineBlock, 256
RefineBlock, 256 CondRefineBlock, 256
RefineBlock, 128 CondRefineBlock, 128
RefineBlock, 128 CondRefineBlock, 128
$\times3$ Conv2D, 2 $\times3$ Conv2D, 2
@@ VI-A3 Evaluation Settings
During evaluation, the channel of each device is randomly sampled from the evaluation set. The following settings are adopted unless specified otherwise: $\lambda=0.05$, $K=800$, and the SNR is $10$ dB. For the figures showing the NMSE [^1] The NMSE of channel estimation is defined as $\frac{\sum_{k}\|\mathbf{H}_{k}-\mathbf{H}_{k}^{\mathtt{post}}\|_{F}^{2}}{\sum_{k}\|\mathbf{H}_{k}\|_{F}^{2}}$. performance and detection error probability, each data point is averaged over $500$ and $5000$ Monte Carlo trials, respectively. The evaluation utilizes a single NVIDIA A800 GPU, which is performed on the same machine used for training.
@@ VI-C JADCE Performance
Fig. 6 shows the NMSE in channel estimation versus the SNR. Our method demonstrates a significant improvement over the baselines across all SNRs. The performance gap between our method and BG-TMP highlights the critical importance of an accurate channel prior in message passing. Moreover, the TMP-EM algorithm favors a smaller number of sub-blocks at low SNR, while preferring a larger number of sub-blocks at high SNR. The sensitivity to the number of sub-blocks can pose challenges in hyperparameter tuning. In contrast, STMP-JADCE learns the channel prior directly from data, eliminating the need for hand-crafted modeling and achieving superior adaptability with consistent performance gains over model-based baselines.
@@
In Fig. 7, we present the detection error probability, which sums up missed detections and false alarms, plotted against different SNRs. The BG-TMP algorithm often fails to converge and is not included. Interestingly, although all schemes utilize the same Bernoulli prior for device activity, our method achieves the lowest detection error probability. This improvement can also be attributed to the plug-in of the score-based generative channel prior. The reason is the following. As shown in Fig. 6, the accurate channel prior brings a remarkable gain in channel estimation. Through message-passing iterations, the device activity is detected using this more accurate channel estimate, hence resulting in enhanced detection performance.
@@
In Fig. 8, we plot the achieved NMSE in channel estimation for various numbers of total devices. Targeted at a channel estimation NMSE of 25-25 dB, our method can support $3000$ connected devices, while the best baseline (TMP-EM with $12$ sub-blocks) can accommodate only $750$ devices, leading to a $4\times$ increase in access capacity. Moreover, for NMSE values below 5-5 dB, STMP-JADCE shows the flattest slope in the performance curve. This suggests that, given an equivalent tolerance in NMSE degradation (in dB), our approach can support the largest increase in the number of connected devices. These observations highlight the superiority of the proposed algorithm for achieving massive connectivity.
@@
Figure 11: NMSE in channel estimation versus SNR for cross-CDL evaluation, where $\lambda=0.05$, $K=800$, and $T=20$. The score networks are trained on the CDL-C dataset, while the STMP-JADCE algorithm is evaluated on the distinct CDL-B channel model.
Figure 12: Convergence behaviors of STMP-JADCE under different measurement matrices, where $\lambda=0.05$, $K=1024$, $T=30$, and $\gamma=0.8$.
Figure 13: NMSE in channel estimation versus the number of training samples, where $\lambda=0.05$, $K=800$, $T=30$ and the SNR is $3$ dB.
To examine the robustness of STMP-JADCE under channel model mismatch, we conduct a cross-CDL evaluation in which the score networks are trained on the CDL-C dataset, while the STMP-JADCE algorithm is evaluated on the distinct CDL-B channel model. As shown in Fig. 11, STMP-JADCE maintains strong performance and continues to outperform all baselines across the entire SNR range. This cross-model consistency can be attributed to the randomized generation of angles, delays, and complex path gains in the CDL-C training set, which exposes the score prior to a broad variety of channel realizations, as well as to the inherent generalization capability of the generative score networks.
We further evaluate STMP-JADCE under different measurement matrices without retraining the generative prior. In addition to the DFT-based design, we consider a Hadamard pilot matrix whose entries are binary $\{\pm 1\}$ and mutually orthogonal across rows, providing a real-valued and multiplication-free alternative to complex-valued DFT pilots. As shown in Fig. 12, the algorithm exhibits consistent convergence behavior and comparable NMSE performance across DFT and Hadamard pilots. This consistency arises naturally from the plug-and-play formulation of STMP-JADCE, in which the unconditional score prior operates independently of the pilot structure, allowing seamless adaptation to various pilot designs drawn from suitably random matrix ensembles.
Finally, to assess the data efficiency of the proposed STMP-JADCE, we evaluate its performance under different training set sizes in Fig. 13. The performance rapidly improves as the dataset size increases from $8\times 10^{2}$ to $4\times 10^{3}$, and then quickly saturates beyond $1\times 10^{4}$ samples. This result indicates that STMP-JADCE requires only a moderate amount of data to achieve near-optimal performance, demonstrating strong data efficiency and robustness even when large-scale datasets are unavailable.
@@ VII Conclusion and Discussion
In this paper, we introduced STMP-JADCE for massive connectivity, a model-based deep learning solution that combines message passing with score-based generative models. In each message-passing iteration, our approach maintains both a first-order and a second-order score network to calculate the posterior mean and variance of the channel, respectively. We show by simulation results that STMP-JADCE drastically outperforms all competing algorithms, while achieving a comparable or even faster convergence speed.